Air of 400kg was moving at 200/s pinged of the vein of a windmill every second.
A) At what rate in KW is the energy arriving in the windmill
B) What is the maximum mass of water that could be bumped each second through a vertical height of 5m?
1
Expert's answer
2019-10-16T09:27:52-0400

A)


"P=0.5\frac{dm}{dt}v^2"

"P=0.5(400)200^2=8000000\ W=8000\ kW"

B)

"\Delta M=\frac{P}{gh}\Delta t"

"\Delta M=\frac{8000000}{(5)(10)}1=160000\ kg"